OpenStudy (anonymous):

The sum of two numbers is 90. Their difference is 12. What are the numbers? Select one: a. no solution b. 31 and 59 c. 35 and 47 d. 39 and 51

OpenStudy (alekos):

let the numbers be a an b a + b = 90 and a - b = 12

OpenStudy (alekos):

Now solve

OpenStudy (anonymous):

all u have to do is add two numbers like 31+59=?your answer it has to equal 90 cause its asking to find 2 #'s that equal 90 so check A first and se if u get 90 as a answer

OpenStudy (anonymous):

sorry I meant b so ...31+59=90 so B is ur answer

OpenStudy (anonymous):

actually it is c because the numbers have to have a difference of 12

OpenStudy (anonymous):

if the answer 31+59 and equals 90 doesn't that answer already have a difference of 12

OpenStudy (anonymous):

no it has a difference of 28 @daddysgirl14

OpenStudy (anonymous):

ok then its C ....

OpenStudy (anonymous):

it's 39 and 51

OpenStudy (anonymous):

thank you for trying though(: @daddysgirl14

OpenStudy (anonymous):

thx add me if u want

OpenStudy (anonymous):

@annas we already figured this out it is c not b

OpenStudy (anonymous):

51 - 39 = 12 :3

OpenStudy (anonymous):

hi everyone

OpenStudy (anonymous):

51-39 does equal 12 but im still think its C

OpenStudy (anonymous):

so it could be c or d but I'm going with c

OpenStudy (anonymous):

x+y = 90 x-y = 12 Add the equations 2x = 102 x = 102/2 x = 51 putting x in equation 1 (51) + y = 90 y = 90 - 51 y = 39 so x= 51 , y= 39 51 - 39 = 12

OpenStudy (anonymous):

Please dont guide her wrong. Thank you!

OpenStudy (anonymous):

51 + 39 = 90 too

OpenStudy (anonymous):

i said that at first too